OpenStudy (anonymous):

4(x – 3) = 5(x + 1) Solve the equation above for x

OpenStudy (anonymous):

4(x - 3) = 5(x+1) 4x -4*3 = 5x +5*1 solve for x

OpenStudy (anonymous):

is it 17

OpenStudy (anonymous):

its -17

OpenStudy (anonymous):

okay i see thanks

OpenStudy (anonymous):

anytime :)